Resourceful Corporation is considering the implementation of a JIT inventory system. The company recently analyzed its cycle time to determine the average number of days spent in each activity of its production process. A summary of the analysis is shown below: Production Activity Receiving materials Inspecting materials Storing materials Moving materials into production Setting-up production equipment Cutting materials Assembling materials Painting finished products Packaging finished products Number of Davs 4 4 Resourcefuls value-adding production activiti es include b Resourcefuls total cycle time is c Resourcefuls manufacturing efficiency ratio is d Which activities might bereduced or eliminated should Efficient implement a JIT system? days
Answer #1

a)Cutting materials, Assembling materials, Painting finished products, Packaging finished products.(6+7+4+1 days = 18 days)

b)1+4+9+6+7+6+7+4+1=45 days

c)18/45 = 40%

d)Receiving materials, storing materials, moving materials into production
you cannot eliminate all these but reduce time
like receiving materials on a particular day of the month and keeping enough safety stock
storing materials , like ordering when the safety stock is started to be used
moving materaisl directly into production rather than storing